Group Selection Page HELP Controls Group ITDM Vacuum WP78 FXE SPB HED SQS SXP Sample Environment PSPO XO TS Migrated2Zulip Archived2GitlabPages Archived2PDF MigratedByGroup
SCS Photon Commissioning Team XRO DOC WP74 MID XO_mig XO_Operation SQS Migrated2Zulip DET
Technical integration and coordination
  Technical integration and coordination logbook, Contact: A. Violante, G. Wellenreuther  Not logged in ELOG logo
Message ID: 93     Entry time: 18 Dec 2019, 10:06
Author: Adriano Violante 
Location: General 
Severity: Information 
Subject: Report after tech. Int. meeting 09.30 

MID:

  • SDL tests ongoing. Important motors have been tested. Alignment of the mechanics with the alignment laser ongoing. Will be pumped over Christmas.

 EEE:

  • uTCA puppet migration: Instr. done. Tunnels and balconies probably will be updated at the beginning of next year.
  • SASE2 EPS loop will be updated probably today.
  • SPB/SFX is getting a new CPU.
  • Migration to MC2 motors has been agreed with controls, XRO, XPD. in XTD 6 XPD the upgrade will be done in KW2. For the rest a plan has to be made.

ITDM

  • Meeting this week ITDM/EEE/TS to discuss how to mitigate issues caused by accidents (e.g. cooling water interruption caused by power cut). 

Controls:

  • CY sent a mail several weeks ago to XO and the DH. The topic is emergency procedures during Christmas. No reply. CY announces that in the period 23.12-07.01 the Controls OCDs would not be available.

AoB

  • No dispatch meeting on 06.01. Next meeting will be the Technical Integration one on 08.01.
ELOG V3.1.4-7c3fd00